OpenStudy (anonymous):

If u = <5, -4>, find u · u. 3 9 sqrt(41) 41

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

I dont think there is a V

OpenStudy (astrophysics):

\[<5,-4> \cdot <5, -4> \implies <5 * 5+ (-4) * (-4)>\]

OpenStudy (astrophysics):

That's the dot product

OpenStudy (michele_laino):

we have, by definition: \[u \cdot u = 5 \times 5 + \left( { - 4} \right) \times \left( { - 4} \right) = ...?\]

OpenStudy (anonymous):

I got a 25 and a 16

OpenStudy (anonymous):

and thas a 41

OpenStudy (astrophysics):

Yup